OpenStudy (anonymous):

solve for x x^4-8x^2-9=0

OpenStudy (anonymous):

x=i,-i,3,-3

OpenStudy (anonymous):

thanks how did you find the answers

myininaya (myininaya):

let u=x^2 so u^2=x^4 u^2-8u-9=0 (u+1)(u-9)=0 u=-1 u=9 remember u=x^2 x^2=-1 implies x=plus or minus i x^2=9 imples x=plus or minus 3